Human colour vision is “produced” by the nervous system based on how three different cone receptors interact with photons of light in the eye. These three different types of cones interact with photons of different frequency light, as indicated in the following chart:

Cone Type Range of Light Frequency Detected

S 6.00-7.49 x 1014 s-1

M 4.76-6.62 x 1014 s-1

L 4.28-6.00 x 1014 s-1


What wavelength ranges (and corresponding colours) do the three types of cones

detect?


1
Expert's answer
2022-04-19T03:11:57-0400

"\lambda" = c / v = 3 x 108 m/s / (6.00-7.49 x 1014 s-1) = 5.00-4.00 x 10-7 m = 400-500 nm

"\lambda" = c / v = 3 x 108 m/s / (4.76-6.62 x 1014 s-1) = 6.30-4.53 x 10-7 m = 453-630 nm

"\lambda" = c / v = 3 x 108 m/s / (4.28-6.00 x 1014 s-1) = 7.00-5.00 x 10-7 m = 500-700 nm

  • Violet: 380–450 nm (688–789 THz frequency)
  • Blue: 450–495 nm
  • Green: 495–570 nm
  • Yellow: 570–590 nm
  • Orange: 590–620 nm
  • Red: 620–750 nm (400–484 THz frequency)